A courier service company wishes to estimate the proportion of people in various states that will use its services. Suppose the true proportion is 0.04. If 223 are sampled, what is the probability that the sample proportion will be greater than 0.05? Round your answer to four decimal places.
Answer #1

Here, μ = 0.04, σ = 0.0131 and x = 0.05. We need to compute P(X >= 0.05). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (0.05 - 0.04)/0.0131 = 0.76

Therefore,
P(X >= 0.05) = P(z <= (0.05 - 0.04)/0.0131)
= P(z >= 0.76)
= 1 - 0.7764 = 0.2236